OpenStudy (anonymous):

Solve for the variable. Reduce the answer to lowest terms. 24q = 240

OpenStudy (anonymous):

\[24q=240\] \[q=\frac{240}{24}=10\]

OpenStudy (anonymous):

if you need an additional step it is this \[24q=240\] \[\frac{1}{24}\times 24q=\frac{1}{24}\times 240\] \]1q=\frac{240}{24}\] \[q=10\]

OpenStudy (anonymous):

\[1q=\frac{240}{24}\]